OpenStudy (koury_with_a_k):

Help Please?! A 0.145 kg baseball is thrown horizontally from a platform, so that it starts at 5.2 m above the ground. How much has its kinetic energy increased when it reaches the ground? 16 J 7.4 J 24 J 0 J Can someone please walk me through this, I'd like to know how to solve it.

OpenStudy (anonymous):

Do you know all the formulas? like Kinetic Energy is \(\frac{1}{2}mv^{2}\) and how to find velocity?

OpenStudy (koury_with_a_k):

No. That's why I need someone to walk me through it. Sorry.

OpenStudy (anonymous):

No worries. So I am assuming this is either hs level or like 150?

OpenStudy (anonymous):

So kinetic energy is \[1/2*m*v^2\]

OpenStudy (koury_with_a_k):

Okay, seems easy enough.

OpenStudy (anonymous):

And to find velocity they should give you a few equations. I need to grab the best one. ONe second

OpenStudy (anonymous):

Do you know what acceleration it will be subject to?

OpenStudy (koury_with_a_k):

No...

OpenStudy (anonymous):

Well the ball will fall to the floor. What will take it, or accelerate it in that direction?

OpenStudy (koury_with_a_k):

Gravity?

OpenStudy (anonymous):

They should provide you with the equation \[v(final)^2 = v(inital)^2 + 2a(x(final)-x(initial))\]

OpenStudy (anonymous):

Yes gravity is an acceleration. Do you know its quantity?

OpenStudy (koury_with_a_k):

9.8 something, I believe.

OpenStudy (anonymous):

Yes, 9.81 m/s^2 is the accepted number. Or 32 ft/s^2 if you are using American Units. Which we are using metric

OpenStudy (anonymous):

So knowing that, we want to find the change in KE. So when we look at the KE equation, this will depend on the velocity. Since the mass cannot change. Correct?

OpenStudy (anonymous):

So at first I thought we needed a horizontal velocity for the ball. But, in reality this will not matter since in this case the horizontal velocity will not change, there is no horizontal acceleration. And we are looking only for a change in KE, or a change in Velocity

OpenStudy (anonymous):

So using the longer equation I provided, you can look at it in terms of 1 dimension. The vertical dimension. So look at the velocity initial, vertically?

OpenStudy (koury_with_a_k):

I'm sorry, how do you find initial velocity?

OpenStudy (koury_with_a_k):

I don't have any of my textbooks with me so I can't look it up.

OpenStudy (anonymous):

Ok, well the initial velocity will have 2 dimensions. The horizontal is not needed, since it doesnt change and therefor will not change. The initial velocity in the vertical is relative. so the change in velocity is equal to acceleration times time.

OpenStudy (anonymous):

Since they dont tell you, assume it is 0. And the change in vertical velocity, or velocity final minus initial will be what matters.

OpenStudy (anonymous):

So if velocity initial is 0, we can use algebra to make the longer equation a little shorter. Like the change in X vertical will also be easy to simplify. Since it falls 5.2 m or whatever.

OpenStudy (anonymous):

So now you have acceleration (gravity), velocity initial (0), x final - x initial (change in height), and you want to know velocity final. You have one unknown, and one equation. You can solve for the unknown.

OpenStudy (anonymous):

You would then take your velocity final from the equation, and plug it into the KE equation. Keep in mind that this velocity final is actually the change in velocity of the ball, assuming the ball has a 0 velocity in the vertical direction. Therefore the KE you get with the velocity is in fact the change in KE, not the balls actual KE.

OpenStudy (koury_with_a_k):

Okay, I think I'm missing something. I have... v(final)^2 = (0) + 2(9.8 m/s^2)(5.2 m) 102 m^2/s^2 v(final)^2 = 10 m/s KE = (1/2)(0.145kg)(10m/s) KE = 0.725 This is not one of my answers though.

OpenStudy (anonymous):

v^2 = 100

OpenStudy (anonymous):

its 1/2*m*v^2 = 0.145*102.024/2 = 7.4

OpenStudy (koury_with_a_k):

Oh okay, I see now. Thank you so much. I greatly appreciate it.

OpenStudy (anonymous):

No problem